OpenStudy (anonymous):

Can someone explain to me what factoring is? Because every time my teacher tries my eyes just glare over and I get really confused like the kind where it's like (x-/+ ?)(x-/+ ?)

OpenStudy (anonymous):

x^2+2x+4=(x+2)(x+2)=(x+2)^2..... Happy Studying!!!!!